# Exploring Feedback Generation in Automated Skeletal Movement Assessment: A Comprehensive Overview

### v2

@@ 2 Movement Assessment
Movement assessment, as one of the applications of movement analysis from skeletal videos, presents a relatively challenging task. Unlike action recognition, which typically involves classifying movements from a high-level temporal and skeletal perspective, movement assessment requires recognizing low-level details crucial for performance quality and closely attending to them at specific stages of movement progress [21]. Automatic movement assessment solutions find application across various domains, including rehabilitative medical assessments [46, 22, 17, 36, 29, 41, 5, 48, 6, 50, 34, 13, 31, 28], gymnastics [43, 4, 26, 35], physical skills [15, 25, 49], and Olympic sports [32].
@@
Feedback generation, which holds the potential to accelerate movement learning and, consequently, the rehabilitation process [44, 19, 21, 10], is occasionally addressed by studies focusing on movement assessment [3, 22, 29, 31, 43, 26, 33]. Feedback is sometimes equated with interpretability, which entails understanding why a specific outcome is observed how input changes will alter it [19]. As we discuss in this paper, the concept of feedback generation can be interpreted in multiple ways. Evaluating the quality of generated feedback can be challenging. From a data perspective, annotating training and test samples with their expected feedback may be considerably more complex than solely annotating them with their expected performance scores. Moreover, in certain scenarios, multiple feedback outputs may be deemed correct or partially correct, posing challenges for evaluation. In the subsequent sections of this paper, we delve deeper into feedback generation.
@@ 3 Feedback Generation
Defining ”feedback” can be complex. In the context of movement assessment, it encompasses a wide array of response types, occurring during or after a performance. Feedback may consist of both basic and advanced guidance, conveyed verbally [3, 26, 31], textually [1, 22, 31, 29], or visually [3, 1, 11, 43, 26], and occasionally through physical means [20]. Table 1 provides a summary of feedback generation solutions, based on the criteria proposed in the following sections.

### v3

@@ 2 Movement Assessment
Movement assessment, as one of the applications of movement analysis from skeletal videos, presents a relatively challenging task. Unlike action recognition, which typically involves classifying movements from a high-level temporal and skeletal perspective, movement assessment requires recognizing low-level details crucial for performance quality and closely attending to them at specific stages of movement progress [23]. Automatic movement assessment solutions find application across various domains, including rehabilitative medical assessments [50, 24, 18, 38, 31, 44, 5, 52, 6, 54, 36, 14, 33, 30, 45], gymnastics [47, 4, 28, 37], physical skills [16, 27, 53], and Olympic sports [34].
@@
Feedback generation, which holds the potential to accelerate movement learning and, consequently, the rehabilitation process [48, 20, 23, 11, 8], is occasionally addressed by studies focusing on movement assessment [3, 24, 31, 33, 47, 28, 35]. Feedback is sometimes equated with interpretability, which entails understanding why a specific outcome is observed how input changes will alter it [20]. In other cases, it is referred to as guidance [22]. As we discuss in this paper, the concept of feedback generation can be interpreted in multiple ways. Evaluating the quality of generated feedback can be challenging. From a data perspective, annotating training and test samples with their expected feedback may be considerably more complex than solely annotating them with their expected performance scores. Moreover, in certain scenarios, multiple feedback outputs may be deemed correct or partially correct, posing challenges for evaluation. In the subsequent sections of this paper, we delve deeper into feedback generation.
@@ 3 Feedback Generation
Defining ”feedback” can be complex. In the context of movement assessment, it encompasses a wide array of response types, occurring during or after a performance. Feedback may consist of both basic and advanced guidance, conveyed verbally [3, 28, 33], textually [1, 24, 33, 31, 22], or visually [3, 1, 12, 47, 28, 40], and occasionally through physical means [21]. Table 1 provides a summary of feedback generation solutions, based on the criteria proposed in the following sections.
@@ 4 Evaluating Generated Feedback
[40] 2023 Physical Rehabilitation Visual Unbound Online Local Qualitative
[22] 2023 Physical Rehabilitation Text Unbound (order=1) Offline N/A Quantitative(Q)